OpenStudy (eric_d):

The expression x^3-4x^2+x+6 and x^3-3x^2+2x+k hv a common factor.Find possible values of k. @ganeshie8

ganeshie8 (ganeshie8):

@DLS

OpenStudy (eric_d):

@DLS

OpenStudy (dls):

why am I being tagged lmfao

OpenStudy (dls):

I ain't even getting notification of being tagged!

OpenStudy (aum):

x^3-4x^2+x+6 Try to factor this. Try a few small values for x such as +1, -1, +2, -2, etc. until you find one factor. x = 1: 1 - 4 + 1 + 6 = 4 x = -1: -1 - 4 -1 + 6 = 0. (x+1) is a factor. Do synthetic division of: (x^3-4x^2+x+6) / (x+1). Find the quotient. Factor the quadratic and find the other two factors.

OpenStudy (eric_d):

factors of 6... (x+1)(x+2)(x-3) something like this?

OpenStudy (dls):

Alright,I'll give it a shot =] The expression x^3-4x^2+x+6 and x^3-3x^2+2x+k hv a common factor.Find possible values of k. Lets take up the first equation. x^3-4x^2+x+6=0 We try to find ONE factor by hit and trial to apply remainder factor theorem. We usually try small values like x=0,1,-1. If you observe, "-1" satisfies the equation (i)...so..(x+1) is a factor here. Now we divide the equation by its factor i.e x+1 and we get the required stuff.

OpenStudy (dls):